# Javier Alejo, Antonio F. Galvao, Julian Martinez-Iriarte, and Gabriel Montes-Rojas

### v2

@@ Abstract
Linear regressions with endogeneity are widely used to estimate causal effects. This paper studies a framework that has two common issues, endogeneity of the regressors, and heteroskedasticity that is allowed to depend on endogenous regressors, i.e., endogenous heteroskedasticity. We show that the presence of such conditional heteroskedasticity in the structural regression renders the two-stages least squares estimator inconsistent. To solve this issue, we propose sufficient conditions together with a control function approach to identify and estimate the causal parameters of interest. We establish the limiting properties of the estimator, say consistency and asymptotic normality, and propose inference procedures. Monte Carlo simulations provide evidence of the finite sample performance of the proposed methods, and evaluate different implementation procedures. We revisit an empirical application about job training to illustrate the methods.
Keywords: Two-stages least squares, instrumental variables, heteroskedasticity, control function.
@@ 1 Introduction
This paper studies linear models where endogeneity and heteroskedasticity are intertwined. At the estimation stage, endogeneity is dealt with instrumental variables (IV) and two-stages least squares (2SLS) procedures.1112SLS methods have been employed in many recent and varying contexts, as for example, among others, estimation of causal effects (Mogstad, Torgovitsky, and Walters (2021)), recursive form of the general simultaneous equations models (Hausman (1983)), local average treatment effects (Imbens and Angrist (1994)), regression discontinuity (Hahn, Todd, and Van der Klaauw (2001)), etc. On the other hand, heteroskedascity is taken care with robust inference procedures. [^2] It has been standard in the literature to account for some form of heteroskedastiticy in the structural model, see for example, the large recent literature on inference with clustering. The seminal contributions to produce inference procedures accounting for heteroskedasticity are White (1980), Pagan and Hall (1983) and Newey and West (1987). More recently, since Arellano (1987), Wooldridge (2003a), and Bertrand, Duflo, and Mullainathan (2004), there has been a large literature on using cluster robust covariance methods to correct size distortions when performing inference. The cluster robust inference has found a wide audience in applied economics and finance research. The cluster robust method has been further developed by Kézdi (2004), Donald and Lang (2007) and Hansen (2007). Cameron and Miller (2015) provide a survey of advancements in this literature. Moreover, papers like Raj, Srivastava, and Ullah (1980) and White (1982) implement heteroskedasticity robust inference for 2SLS models. When heteroskedasticity is caused by an endogenous regressor, by capturing heterogeneity in the main effects of interest, the 2SLS estimator is inconsistent because the required exogeneity condition does not hold. The literature studying econometric models allowing for endogenous heteroskedasticty simultaneously is very sparse (see, e.g., Florens, Heckman, Meghir, and Vytlacil (2008), Chen and Khan (2014) and Abrevaya and Xu (2023)) and this paper specifically focus on this.
@@
We provide alternative methods for identification, estimation, and inference for the parameter of interest in linear triangular models with endogeneity and endogenous heteroskedasticity. We study identification using a control function (CF) approach in the presence of heteroskedasticity in both the structural and first-stage equations. [^3] There is a large literature on the CF approach. See, e.g., among many others, Newey, Powell, and Vella (1999) and Kim and Petrin (2022). Of particular interest are linear triangular systems is an issue of interest for econometricians together with the CF approach. Different alternatives have been studied in the literature. See the survey by Blundell and Powell (2008) and Imbens and Newey (2009) for nonparametric and semiparametric models. We first show that a general nonparametric CF identification cannot be achieved because of the skedastic terms. We then build on Florens, Heckman, Meghir, and Vytlacil (2008) and provide conditions to solve this issue. The relevant restrictions for identification are a CF condition, together with the use of a flexible parametrization of the skedastic function in the structural equation. Given the available CF from the first-stage, and the structure of the skedastic function for the structural equation, a simple ordinary least squares (OLS) can be used to identify the parameter of interest in the second-stage.
@@
Given the identification result, estimation is simple and can be implemented empirically in a two-step procedure. First, the skedastic function in the first step can be estimated using flexible scale function methods suggested in Romano and Wolf (2017) and Wooldridge (2010). [^4] The parametric model in the first-stage is mainly used due to the curse of dimensionality problem, since most empirical work uses multiple control variables. In a nonparemetric case, estimation could rely on methods suggested by Jin, Su, and Xiao (2015) and Linton and Xiao (2019). Given the estimate of the skedastic function, the CF can be computed as the normalized errors from the first-stage regression. [^5] We note that under heteroskedasticity in the first-stage regression, by assuming the availability of a valid exogenous IV, the CF can be identified by using a flexible parametric model for the skedastic function. By normalizing the variance of the error term in the first-stage equation, a valid CF variable can be constructed. In the second step, one performs a simple linear regression with a CF approach. Empirically, this step uses an OLS regression of the dependent variable on the endogenous, exogenous, and CF variables augmented with their interacting terms. We establish the limiting statistical properties of the proposed estimator. Mild sufficient conditions are provided for the two-step estimator to have desired asymptotic properties, namely, consistency and asymptotic normality. In addition, we develop statistical inference procedures by estimating the asymptotic variance-covariance matrix with the generated regressors.
@@
We also evaluate the finite sample performance of the proposed methods using Monte Carlo exercises. Numerical simulations confirm that the presence of heteroskedasticity in the structural equation induces bias in the 2SLS estimator. Moreover, heteroskedasticity in the first-stage equation may amplify the bias. The proposed CF method is able to produce unbiased estimates. In addition, results improve when the sample size increases.
@@
Related literature. The literature on triangular systems of equations with a continuous treatment is large, see, e.g., among many others, Newey, Powell, and Vella (1999), Chesher (2003), Newey and Powell (2003), Ma and Koenker (2006), Lee (2007), Imbens and Newey (2009), D’Haultfœuille and Février (2015), Torgovitsky (2015), Han (2020), Kim and Petrin (2022), and Pereda-Fernández (2023). There is also a large literature on heterogeneous causal effects see, e.g., Imbens and Angrist (1994), Wooldridge (1997), Heckman, Smith, and Clements (1997), Wooldridge (2003b), Heckman and Vytlacil (2005), Heckman and Vytlacil (2007), Carlson (2023), and Chen, Khan, and Tang (2024).
@@
The literature considering the effects of endogenous heteroskedasticity is more restricted. When treatment effects are heterogeneous among observationally identical individuals, that is, in the presence of heteroskedasticity in the structural equation, causal inference for policy evaluation is known to be more difficult. Florens, Heckman, Meghir, and Vytlacil (2008) use the control function approach to identify the average treatment effects and the effect of treatment on the treated in models with a continuous endogenous regressor whose impact is heterogeneous. Abrevaya and Xu (2023) show that when a binary treatment is endogenous and there is endogenous heteroskedasticity (EH), that is, the endogenous treatment indicator also affects the scale of the outcome variable, then standard methods for estimation of average treatment effects fail. In particular, they consider the standard IV estimator with binary treatment and show that it is an inconsistent estimator for the average treatment effect. After nonparametric identification is established, they provide closed-form estimators are provided under a linear specification for the mean and variance treatment effect, as well as the average treatment effect on the treated. In the context of local average treatment effects, Chen and Khan (2014) discuss identification and estimation of the heteroskedasticity term under different treatment statuses. We show that these are particular cases of a general model, since if the endogenous regressor is a dummy variable, then the first-stage in a 2SLS model has heteroskedasticity by construction. Then EH-IV within this framework produces the type of inconsistency we are concerned of. Note that this may occur even if the IV is randomly generated, as frequent in experimental frameworks.
@@
This paper is organized as follows. Section 2 presents the main model, reviews the 2SLS methods, and its inconsistency under endogenous heteroskedasticity. In Section 3 we propose a CF approach, discuss identification, estimation, and inference procedures. A Monte Carlo study is provided in Section 5. Section 6 illustrates the methods with an empirical application. Finally, Section 7 concludes. Mathematical proofs are relegated to the Appendix.
2 Main Model
This section introduces the model we study in this paper, which encompasses endogeneity and heteroskedasticity. We then review the two-stages least squares (2SLS) strategy for estimating the parameter of interest. We show that the 2SLS is inconsistent in the presence of endogenous heteroskedasticity.
2.1 Location-scale models with endogeneity
We consider the following model with endogenous heteroskedasticity:
@@ 2.1 Location-scale models with endogeneity
$D$ $=Z^{\prime}\pi_{1}+X^{\prime}\pi_{2}+h(X,Z)V.$ (2)
In this model, $Y$ is the response variable, $D$ is an endogenous variable, and the main parameter of interest is $\alpha_{1}$, which is real-valued. [^6] We present results for a scalar endogenous variable for simplicity. The results are simple to extend to the general case of multiple endogenous variables. The vector of exogenous variables $X$ is $(p_{x}-1)$-dimensional (including a constant), and the instrumental variable $Z$ is $p_{z}$-dimensional. The variables $\varepsilon$ and $V$ are the innovation terms. Here, the correlation between unobservable innovations $\varepsilon$ and $V$ renders $D$ to be potentially endogenous. The heteroskedasticity inducing functions are $g:\mathbb{R}^{p_{x}}\mapsto\mathbb{R}_{>0}$ and $h:\mathbb{R}^{p_{z}+p_{x}-1}\mapsto\mathbb{R}_{>0}$. The functions $g$ and $h$ are unknown and when $g\equiv 1$ ($h\equiv 1$) there is homoskedasticity in the structural equation (first-stage equation). To complete the model, we assume that $E[\varepsilon|X,Z]=0$ and $E[V|X,Z]=0$. But we will discuss the conditional expectation restrictions on these random variables in details later.
The model in (1)–(2) is a location-scale model that allows for endogeneity and heteroskedasticity. This is a flexible general structural model that can be employed in a variety of contexts. It is typically used to evaluate heterogeneity in treatment effects as in Florens, Heckman, Meghir, and Vytlacil (2008) and Abrevaya and Xu (2023). It is also used to model the impact of heterogeneity on the conditional distribution of the outcome variable via conditional quantiles as in Ma and Koenker (2006) and Lee (2007).
Next, we review the properties of the 2SLS estimator, which is the most common method used to estimate parameters under endogeneity. We show that in the presence of the skedastic functions renders that the 2SLS identification strategy to fail under the usual assumptions in the literature, and consequently the estimator is inconsistent.
2.2 Review of 2SLS and intuition
Before we proceed to a general result on the property of the 2SLS under endogenous heteroskedasticity, we briefly review the 2SLS identification of the model in (1)–(2), in the context of estimation of treatment effects. Consider, first, the following simplified structural model under endogeneity
$Y$ $=D\alpha_{1}+\varepsilon,$ (3)
where $Y$ is the response, $D$ is the treatment, $\alpha_{1}\in\mathbb{R}$ is the parameter of interest, $\varepsilon$ is the unobserved innovation with $E[\varepsilon]=0$ but $E[\varepsilon|D]\neq 0.$ Suppose that a valid instrumental variable (IV), $Z\in\mathbb{R}^{p_{z}}$, is available, such that it satisfies the standard moment conditions $E[\varepsilon|Z]=0$, and $Z$ is correlated with $D$ [^7] We note that the weaker condition $E[Z\varepsilon]=0$ also suffices. Let $\Pi:=E[ZZ^{\prime}]^{-1}E[ZD]$, the 2SLS uses the one dimensional instrument $\Pi^{\prime}Z$. Then, $\alpha_{1}$ is identified by taking expectations on $\Pi^{\prime}ZY=\Pi^{\prime}ZD\alpha_{1}+\Pi^{\prime}Z\varepsilon$. Under the usual rank conditions, this yields identification of $\alpha_{1}$ by $(\Pi^{\prime}E[ZD])^{-1}\Pi^{\prime}E[ZY]$.
Now suppose that we augment the model in equation (3) with the function $g(\cdot)$ and use the following location-scale model
$Y$ $=D\alpha_{1}+g(D)\varepsilon,$ (4)
and we maintain the same exogeneity condition for the IV: $E[\varepsilon|Z]=0$. However, we note that since $E[\varepsilon|Z]=0$ does not imply $E[Zg(D)\varepsilon]=0$, then the previous identification strategy fails. [^8] This stands in contrast with the case where $D$ is exogenous. In this case, if $E[\varepsilon|D]=0$, then we still have $E[g(D)\varepsilon|D]=0$, and the (heteroskedastic-)OLS strategy still works. To see this, suppose $g(D)=D$, and $D=Z\pi+V$. Then $E[Zg(D)\varepsilon]=E[V\varepsilon]\neq 0$ which is not 00.999Note that the stronger assumption of $Z\perp\varepsilon$ also is not enough. Moreover, equation (4) is a general set-up for heteroskedastic models, provided that for any model $Y=D\alpha_{1}+U,\ Var(U|D)=g(D)^{2}$ we can define the homoskedastic error term $\varepsilon=U/g(D)$.
In the treatment effects literature, the parameter $\alpha_{1}$ can be interpreted as a causal effect. Also, the heteroskedascity function $g(\cdot)$ in equation (4) is used to model heterogeneous treatment effects. It should be noted, however, that our interest lies in estimating $\alpha_{1}$, which can be interpreted as the average treatment effect (ATE), but not as the local ATE (LATE), which is in fact the corresponding parameter that would be obtained if 2SLS were implemented. Indeed, consider the simple case where $Y(d)$ is the potential outcome associated with $D=d$. Then, equation (4) is $Y(d)=d\alpha_{1}+g(d)\varepsilon,$ and so the difference $Y(d^{\prime})-Y(d)$ has a common component, $(d^{\prime}-d)\alpha_{1}$, as well as an heterogeneity component given by $(g(d^{\prime})-g(d))\varepsilon$, which is random through the term $\varepsilon$ (see Florens, Heckman, Meghir, and Vytlacil (2008) and Abrevaya and Xu (2023)). See Appendix A.1 for further details on the interpretation of the parameter $\alpha_{1}$.
This paper studies this case of endogenous treatment effects under heterogeneity. Next we show that the failure of 2SLS strategy is due to the endogenous heteroskedasticity, as opposed to when the instruments are intrinsically invalid in the sense of being correlated with unobservables. Later, we will propose to use identification, estimation, and inference based on a control function approach.
@@ 2.2 Review of 2SLS and intuition
The failure of IV to identify $\alpha_{1}$ in (4) has been pointed out by Abrevaya and Xu (2023) in the case of a binary $D$. When a binary IV is available, they offer a detailed study of the bias and propose a novel identification and estimation strategy. They refer to their framework as Endogenous Heteroskedasticity IV (EHIV). Moreover, Florens, Heckman, Meghir, and Vytlacil (2008) note that in a model similar to (4), with a continuous $D$, the usual requirements on an IV are not enough. They also offer an alternative identification and estimation approach.
To motivate the approach in this paper, we analyze the following simple example, where we illustrate the impact of the endogenous heteroskedasticity on the 2SLS strategy.
@@
Consider the following simple linear case of model (1)–(2) with $g(D)=(1+\delta D)$. To focus on the implications of the endogenous heteroskedasticity in the structural equation, we consider a general first-stage equation. Thus, the model is given by
@@
The exogeneity assumptions on $Z$ do not imply that $Cov[Z,D\varepsilon]=0$. Therefore, $\alpha_{1}$ is not identified when there is heteroskedasticity in the structural equation, i.e., $\delta\neq 0.$ Note that this problem persists even if $Z\perp\!\!\!\perp(\varepsilon,V)$ as in experimental settings.
Figure 1: Bias in 2SLS. In orange the CF estimates, in blue the 2SLS estimates.
The result of Example 1 is further illustrated numerically in Figure 1, which contains, in blue, the empirical distribution of the 2SLS estimator, for a case where $\alpha=1$. The (asymptotic) bias is quite noticeable. As a preview of our results, our proposed estimators, correctly centered at $\alpha=1$, is shown in orange. [^10] More details about the simulation can be found on Section 5.
Next, we study the inconsistency of the 2SLS in more detail. In particular, we also assess the impact of heteroskedasticity in the first-stage equation.
2.3 Inconsistency of the 2SLS estimator
Next, we provide a formal result deriving the asymptotic bias in the 2SLS estimator. We maintain the following assumptions.
@@ 2.3 Inconsistency of the 2SLS estimator
The next result compares the probability limit of the 2SLS estimator to the target parameter $\alpha_{1}$. We assume that the data is generated according to the model in (1)–(2), and a researcher who is interested in $\alpha_{1}$, instruments $D$ with $Z$ is a 2SLS regression of $Y$ of $D$ and $X$ (which includes a constant). In order to deal with the presence of the additional controls $X$, we use a population version of the Frisch-Waugh-Lovell (FWL) theorem using linear projections. [^11] See Section A.2 in the Appendix for a brief review of linear projections.
@@
Consider the model in (1)–(2). Let $\overline{\alpha}_{1}$ be the probability limit of the 2SLS estimator of $\alpha_{1}$ of a 2SLS regression of $Y$ on $D$ and $X$, with $D$ instrumented by $Z$. Under Assumptions 1 and 2, we have
$\overline{\alpha}_{1}=\alpha_{1}+\Sigma_{h}^{-1}\pi_{1}^{\prime}E[\overline{Z}g(Z^{\prime}\pi_{1}+X^{\prime}\pi_{2}+h(X,Z)V,X)\varepsilon],$ (5)
where $\Sigma_{h}:=\pi_{1}^{\prime}E[\overline{Z}\,\overline{Z}^{\prime}]\pi_{1}$, and $\overline{Z}^{\prime}:=Z^{\prime}-X^{\prime}E[XX^{\prime}]^{-1}E[XZ^{\prime}]$.
We provide some further remarks to understand the intuition and nature of the result in Lemma 1.
  1. 1.
    The asymptotic bias of the 2SLS estimator in equation (5) can be succinctly written as $\Sigma_{h}^{-1}\pi_{1}^{\prime}E[\overline{Z}g(D,X)\varepsilon]$, but this might obscure the fact that $h(\cdot)$ is also playing a role through $D$.
  2. 2.
    If $g(\cdot)\equiv 1$, i.e., there is no heteroskedasticity in the structural equation, then there is no asymptotic bias in the 2SLS estimator because $E[\overline{Z}{g(D,X)\varepsilon}]=E[\overline{Z}{\varepsilon}]=0$ by Assumption 1. The lack of bias is regardless of $h(\cdot)$. For a general $g(\cdot)$, the form of $h(\cdot)$ may matter.
  3. 3.
    If $h(\cdot)\equiv 1$, i.e., there is no heteroskedasticity in the first-stage, then the asymptotic bias depends only on the form of $g(\cdot)$.
  4. 4.
@@
    Even in this simple case, the sign of the bias is not totally obvious.
  5. 5.
@@
  6. 6.
    In this paper, we focus on the 2SLS estimator. Nevertheless, we highlight the same result is valid for a GMM estimator. The main intuition is that, under endogenous heteroskedasticity, the moment condition used for estimation would be misspecified trough the function $g(\cdot)$.

### v3

# Julian Martinez-Iriarte, and Gabriel Montes-Rojas

@@ Abstract
Linear regressions with endogeneity are widely used to estimate causal effects. This paper studies a framework that involves two common issues: endogeneity of the regressors and heteroskedasticity that depends on endogenous regressors–i.e., endogenous heteroskedasticity. We show that the presence of endogenous heteroskedasticity in the structural regression renders the two-stage least squares estimator inconsistent. To address this issue, we propose sufficient conditions and a control function approach to identify and estimate the causal parameters of interest. We establish the limiting properties of the estimator–namely, consistency and asymptotic normality–and propose inference procedures. Monte Carlo simulations provide evidence on the finite-sample performance of the proposed methods and evaluate different implementation strategies. We revisit an empirical application on job training to illustrate the methods.
Keywords: Two-stage least squares, instrumental variables, heteroskedasticity, control function.
@@ 1 Introduction
This paper studies linear models with endogeneity and heteroskedasticity. At the estimation stage, endogeneity is typically addressed using instrumental variables (IV) and two-stage least squares (2SLS) procedures. Heteroskedasticity, on the other hand, is usually handled with robust inference techniques. However, when heteroskedasticity is caused by an endogenous regressor, a situation referred to as endogenous heteroskedasticity, the 2SLS estimator becomes inconsistent because the necessary exogeneity condition fails. The literature on econometric models that allow for endogenous heteroskedasticity is very limited (see, e.g., Florens, Heckman, Meghir, and Vytlacil (2008), Chen and Khan (2014), and Abrevaya and Xu (2023)), and this paper focuses specifically on that issue.
@@
We propose alternative methods for identification, estimation, and inference for the parameter of interest in linear triangular models with endogeneity and endogenous heteroskedasticity. We study identification using a control function (CF) approach in the presence of heteroskedasticity in both the structural and first-stage equations. We first show that general nonparametric CF identification fails due to the presence of skedastic terms. We then build on Florens, Heckman, Meghir, and Vytlacil (2008) and provide conditions that address this challenge. The key restrictions for identification are a CF condition, along with a flexible parametrization of the skedastic function in the structural equation. Given a CF constructed from the first stage and the assumed structure of the skedastic function, the parameter of interest in the second stage can be identified using ordinary least squares regression.
@@
Estimation is based on the identification result and it is implemented empirically in a two-step procedure. First, the skedastic function in the first step can be estimated using flexible scale function methods suggested in Romano and Wolf (2017) and Wooldridge (2010). [^1] The parametric model in the first-stage is mainly used due to the curse of dimensionality problem, since most empirical work uses multiple control variables. In a nonparemetric case, estimation could rely on methods suggested by Jin, Su, and Xiao (2015) and Linton and Xiao (2019). Given the estimate of the skedastic function, the CF can be computed as the normalized errors from the first-stage regression. [^2] We note that under heteroskedasticity in the first-stage regression, by assuming the availability of a valid exogenous IV, the CF can be identified by using a flexible parametric model for the skedastic function. By normalizing the variance of the error term in the first-stage equation, a valid CF variable can be constructed. In the second step, one performs a simple linear regression with a CF approach. Empirically, this step uses an OLS regression of the dependent variable on the endogenous, exogenous, and CF variables augmented with their interacting terms. We establish the limiting statistical properties of the proposed estimator. Mild sufficient conditions are provided for the two-step estimator to have desired asymptotic properties, namely, consistency and asymptotic normality. In addition, we develop statistical inference procedures by estimating the asymptotic variance-covariance matrix with the generated regressors.
@@
We also evaluate the finite sample performance of the proposed methods using Monte Carlo exercises. Numerical simulations confirm that the presence of heteroskedasticity in the structural equation induces bias in the 2SLS estimator. Moreover, heteroskedasticity in the first-stage equation may amplify the bias. The proposed CF method is able to produce estimates that concentrate around the true value of the target parameter. In addition, results improve when the sample size increases.
@@
Literature review. There is a large literature on the CF approach. See, e.g., among many others, Newey, Powell, and Vella (1999), Kim and Petrin (2022), and Han and Kaido (2025). Of particular interest are linear triangular systems is an issue of interest for econometricians together with the CF approach. Different alternatives have been studied in the literature. See the survey by Blundell and Powell (2008) and Imbens and Newey (2009) for nonparametric and semiparametric models.
@@
The literature considering the effects of endogenous heteroskedasticity is more restricted. When treatment effects are heterogeneous among observationally identical individuals, that is, in the presence of heteroskedasticity in the structural equation, causal inference for policy evaluation is known to be more difficult. Florens, Heckman, Meghir, and Vytlacil (2008) use the control function approach to identify the average treatment effects and the effect of treatment on the treated in models with a continuous endogenous regressor whose impact is heterogeneous. Abrevaya and Xu (2023) show that when a binary treatment is endogenous and there is endogenous heteroskedasticity (EH), that is, the endogenous treatment indicator also affects the scale of the outcome variable, then standard methods for estimation of average treatment effects fail. In particular, they consider the standard IV estimator with binary treatment and show that it is an inconsistent estimator for the average treatment effect. After nonparametric identification is established, they provide closed-form estimators are provided under a linear specification for the mean and variance treatment effect, as well as the average treatment effect on the treated. In the context of local average treatment effects, Chen and Khan (2014) discuss identification and estimation of the heteroskedasticity term under different treatment statuses. As opposed to the papers listed above, we do not require particular support restrictions for the endogenous variables nor the instrumental variables. That is, we do not require continuity as in Florens, Heckman, Meghir, and Vytlacil (2008), nor discreteness as in Abrevaya and Xu (2023). [^3] Heteroskedasticity can sometimes aid in addressing endogeneity problems rather than exacerbating them, as shown in Rigobon (2003), Klein and Vella (2010), and Lewbell (2011). However, this is not the case considered in the present study.
@@
This paper is organized as follows. Section 2 shows the inconsistency of 2SLS under endogenous heteroskedasticity. In Section 3 we propose a CF approach and discuss identification. Estimation and inference procedures are detailed in Section 4. A Monte Carlo study is provided in Section 5. Section 6 illustrates the methods with an empirical application. Finally, Section 7 concludes. Mathematical proofs are relegated to the Appendix.
2 Inconsistency of the 2SLS estimator
We consider the following model:
@@ 2 Inconsistency of the 2SLS estimator
$D$ $=m(X,Z,V).$ (2)
where the main parameter of interest is the scalar $\alpha_{1}$. The vector of exogenous variables $X$ is $(p_{x}-1)$-dimensional (including a constant), and the instrumental variable (IV) $Z$ is $p_{z}$-dimensional. The variables $\varepsilon$ and $V$ are unobservable and their correlation makes $D$ potentially endogenous.
[…]
@@
The failure of 2SLS to identify $\alpha_{1}$ in the presence of endogenous heteroskedasticity was highlighted by Abrevaya and Xu (2023) in the case of a binary $D$. When a binary IV is available, they offer a detailed analysis of the bias and propose a novel identification and estimation strategy. In a similar context with a continuous $D$, Florens, Heckman, Meghir, and Vytlacil (2008) argue that the usual IV requirements are not sufficient for identification.
The following example illustrates the bias in the 2SLS procedure in a model described by (1)–(2).
@@
Consider the following simple linear case of model (1)–(2) with $g(D)=(1+\delta D)$:
@@
The exogeneity assumptions on $Z$ do not imply that $Cov[Z,D\varepsilon]=0$. Therefore, $\alpha_{1}$ is not identified when there is heteroskedasticity in the structural equation, i.e., $\delta\neq 0.$
The next example shows that this problem persists even if $Z\perp\!\!\!\perp(\varepsilon,V)$ as in experimental settings. To get a closed-form expression we assume a linear first stage.
Example 2.
Consider a special case of Example 1:
$Y$ $=D\alpha_{1}+(1+D\delta)\varepsilon$
$D$ $=Z\pi_{1}+(1+Z\gamma)V,$
The 2SLS estimator is the sample counterpart of $\overline{\alpha}_{1}:=\frac{E[ZY]}{E[ZD]}=\alpha_{1}+\delta\frac{E[ZV\varepsilon]}{\pi_{1}E[Z^{2}]}+\delta\gamma\frac{E[Z^{2}V\varepsilon]}{\pi_{1}E[Z^{2}]}.$ Under the stronger independence assumption: $Z\perp\!\!\!\perp(V,\varepsilon)$, the expression simplifies to
$\overline{\alpha}_{1}$ $=\alpha_{1}+\delta\gamma\frac{E[V\varepsilon]}{\pi_{1}}.$
In this case, the joint heteroskedasticity together with the endogeneity (in the form of $E[V\varepsilon]\neq 0$) is inducing the bias. Thus, in this case, one could have a situation where there is heterogeneity in the structural equation with $\delta\neq 0$, but if there is no heteroskedasticity in the first-stage, $\gamma=0$, the 2SLS is able to correctly identify the parameter of interest.
Next, we provide a formal result deriving the asymptotic bias in the 2SLS estimator under the endogenous heteroskedasticity. This not a novel results per se, but we specify it to our particular setting. We maintain the following assumptions.
@@
The next result compares the probability limit of the 2SLS estimator, denoted by $\overline{\alpha}_{1}$, to the target parameter $\alpha_{1}$. We assume that the data is generated according to the model in (1)–(2), and a researcher who is interested in $\alpha_{1}$, instruments $D$ with $Z$ in a 2SLS regression of $Y$ of $D$ and $X$ (which includes a constant). In order to deal with the presence of the additional controls $X$, we use a population version of the Frisch-Waugh-Lovell (FWL) theorem using linear projections. [^5] See Section A.2 in the Appendix for a brief review of linear projections.
@@
Consider the model in (1)–(2). Let $\overline{\alpha}_{1}$ be the probability limit of the 2SLS estimator of $\alpha_{1}$ of a 2SLS regression of $Y$ on $D$ and $X$, with $D$ instrumented by $Z$ and $X$. Under Assumptions 1 and 2, we have
$\overline{\alpha}_{1}=\alpha_{1}+\Sigma_{Z}^{-1}\pi_{1}^{\prime}E[\overline{Z}g(m(X,Z,V),X)\varepsilon],$ (3)
where $\Sigma_{Z}:=\pi_{1}^{\prime}E[\overline{Z}\,\overline{Z}^{\prime}]\pi_{1}$, and $\overline{Z}^{\prime}:=Z^{\prime}-X^{\prime}E[XX^{\prime}]^{-1}E[XZ^{\prime}]$.
We provide some remarks to further understand the intuition and nature of the result in Lemma 1.
Remark 1.
If $g(\cdot)\equiv 1$, i.e., there is no heteroskedasticity in the structural equation, then there is no asymptotic bias in the 2SLS estimator because $E[\overline{Z}{g(D,X)\varepsilon}]=E[\overline{Z}{\varepsilon}]=0$ by Assumption 1.
Remark 2.
@@
Even in this simple case, the sign of the bias is not obvious.
Remark 3.
@@
Remark 4.
In this paper, we focus on the 2SLS estimator. Nevertheless, we highlight the same result is valid for a GMM estimator. The main intuition is that, under endogenous heteroskedasticity, the moment condition used for estimation would be misspecified through the function $g(\cdot)$.